# Hopf orbits and the first ECH capacity

### v2

@@ 2.5. The main compactness argument
Fix $\delta<0$ in the interval between eigenvalues of the asymptotic operator $L_{\gamma_{+},\widetilde{J}}$ satisfying $\operatorname{wind}_{\tau}=1$ and $\operatorname{wind}_{\tau}=2$. This is possible since it follows from the definition of fast plane that ${\rm CZ}^{0}_{\tau}(\gamma_{+})\geq 3$. Note that $\alpha^{<\delta}_{\tau}(\gamma_{+})=1$ and ${\rm CZ}^{\delta}_{\tau}(\gamma_{+})=3$. Following [props3], one can build a Fredholm theory with weight $\delta$ based on sections of the normal bundle. In more detail, let $\widetilde{u}=(a,u)$ represent an element of $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$. Using the asymptotic behaviour from Theorem 2.4 and the fact that $\widetilde{u}$ is assumed to be an embedding (an immersion would suffice), one can consider a $\bar{J}$-invariant normal bundle $N$ of $\widetilde{u}$ that agrees with $u^{*}\xi$ near the puncture $\infty$. Then one looks at sections of the normal bundle that decay faster then $e^{\delta s}$ near $\infty$, where a neighbourhood of $\infty$ is parametrised by $(s,t)\simeq e^{2\pi(s+it)}$, $s\gg 1$. Here one can use Sobolev or Hölder spaces. The linearization of the Cauchy-Riemann equations at $\widetilde{u}$ determines a Fredholm operator $D_{\widetilde{u}}$ on this space of sections with index
${\rm ind}_{\delta}(\widetilde{u})={\rm CZ}^{\delta}_{\tau}(\gamma_{+})-1=3-1=2.$
@@
We claim that automatic transversality holds, i.e. this Fredholm operator is always surjective, without any genericity assumption on the almost complex structure. To see this, note that a symplectic trivialization of the normal bundle can always be homotoped to a trivialization that extends over a $d\lambda$-symplectic trivialization $\tau_{N}$ of $x(T\cdot)^{*}\xi$. The decomposition of complex vector bundles $\widetilde{u}^{*}TW=T\mathbb{C}\oplus N$ implies that a section of $x(T\cdot)^{*}\xi$ with winding number equal to 00 with respect to $\tau_{N}$ will have winding number equal to $+1$ with respect to $\tau$. In particular, $\alpha^{<\delta}_{\tau_{N}}(\gamma_{+})=\alpha^{<\delta}_{\tau}(\gamma_{+})-1=1-1=0$. A section $\zeta$ in the kernel of $D_{\widetilde{u}}$ satisfies a Cauchy-Riemann type equation. Hence either $\zeta\equiv 0$, or its zeros are isolated and count positively to the total algebraic count of zeros. Moreover, if $\zeta\not\equiv 0$ then its asymptotic behaviour at $\infty$ is governed by an eigenvector $e(t)$ of the asymptotic operator $L_{\gamma_{+},\widetilde{J}}$ associated to an eigenvalue $\nu<0$ in a similar manner to what is described by Theorem 2.4: $\zeta(s,t)\sim e^{\nu s}(e(t)+\varepsilon(s,t))$, $\sup_{t}|\varepsilon(s,t)|\to 0$ as $s\to+\infty$. One can justify such asymptotic behaviour using [sie_CPAM, Theorem A.1]. Hence $\zeta$ has only finitely many zeros, and by the argument principle the total count of zeros of $\zeta$ is nonnegative and equal to $\operatorname{wind}_{\tau_{N}}(\nu)$. The exponential decay $e^{\delta s}|\zeta(s,t)|\to 0$ forces $\nu<\delta$. Thus $\operatorname{wind}_{\tau_{N}}(\nu)\leq\alpha^{<\delta}_{\tau_{N}}(\gamma_{+})=0$, and we conclude that $\zeta$ never vanishes if it does not vanish identically. If $D_{\widetilde{u}}$ is not surjective then its kernel has dimension $\geq 3$ since the index is $2$, so we would find three linearly independent sections in the kernel, but this is impossible: a linear combination of them would vanish at some point because the normal bundle has rank two. This proves the desired surjectivity.
@@
The automatic transversality proved above endows $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ with the structure of a two-dimensional manifold. Note that since the Fredholm theory is modelled on a space of sections with exponential decay faster than $\delta$, the asymptotic eigenvalue of the obtained nearby planes is always $<\delta$, implying that $\operatorname{wind}_{\infty}\leq\alpha^{<\delta}_{\tau}=1$ for the nearby planes.
@@
In preparation for compactness arguments below, we note that one can use results on cylinders of small area such as those proven in [small_area], together with the assumed nondegeneracy of $\gamma_{+}$, to conclude that the topology $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ inherits from the functional analytic set-up agrees with the topology induced by $C^{\infty}_{\rm loc}$-convergence. This is the property of “completeness” in [props3].
@@
Suppose now that $\widetilde{u}$ is a embedded fast plane in $\overline{W}$ asymptotic to $\gamma_{+}$. Let $\mathcal{Y}\subset\mathcal{M}^{\rm fast}(\gamma,\bar{J})$ denote the connected component of $[\widetilde{u}]$. For each $L\geq 0$ consider the compact subset of $\overline{W}$ defined by
@@
$\mathcal{Y}_{L}=\{[\widetilde{v}]\in\mathcal{Y}\mid\widetilde{v}(\mathbb{C})\cap E_{L}\neq\emptyset,\ \widetilde{v}(\mathbb{C})\cap\left((-\infty,L)\times Y^{-}\right)=\emptyset\}.$
@@
Let $[\widetilde{v}_{n}]\in\mathcal{Y}_{L}$ be an arbitrary sequence. By the SFT compactness theorem [sft_comp] we find a subsequence, still denoted by $\widetilde{v}_{n}$, that SFT converges to a holomorphic building ${\bf u}$. The bottom level of the building is a (possibly nodal) curve in $(\overline{W},\bar{J})$, but in principle there might be more levels in (×$(\mathbb{R}\times Y^{+},\widetilde{J})$. Any level in ×$\mathbb{R}\times Y^{+}$ must have at least one irreducible component which is not a trivial cylinder ×$\mathbb{R}\times\gamma$ where $\gamma$ is a Reeb orbit.
We first prove that ${\bf u}$ has only the bottom level. Suppose to get a contradiction that ${\bf u}$ has levels in ×$\mathbb{R}\times Y^{+}$. Then the top level is a finite energy punctured sphere
$\widetilde{u}_{+}=(a_{+},u_{+}):(\mathbb{C}\setminus\Gamma,i)\longrightarrow(\mathbb{R}\times Y^{+},\widetilde{J})$
@@
Note that there is a uniformity in neighbourhoods of $\infty$ encoded in property (i). Write $(s,t)$ instead of $e^{2\pi(s+it)}$. Theorem 2.4 provides $s_{0}>0$ such that $\pi_{\lambda}(\partial_{s}u_{+})$ does not vanish on $[s_{0},+\infty)\times\mathbb{R}/\mathbb{Z}$. In particular the winding number $\operatorname{wind}_{\tau}(\pi_{\lambda}(\partial_{s}u_{+}(s_{0},\cdot)))$ of $t\mapsto\pi_{\lambda}(\partial_{s}u_{+}(s_{0},t))$ in the global frame $\tau$ is equal to $\operatorname{wind}_{\infty}(\widetilde{u},\infty,\tau)$. By (i) we can assume, up to making $s_{0}$ larger, that $\pi_{\lambda}(\partial_{s}w_{n})$ is defined on the set $[s_{0},+\infty)\times\mathbb{R}/\mathbb{Z}$. Use $\tau$ to represent $\pi_{\lambda}(\partial_{s}w_{n})$ as a $\mathbb{C}$-valued function $\zeta_{n}(s,t)$. This function satisfies a Cauchy-Riemann type equation, hence its zeros count positively towards the total algebraic count of zeros. Since $\pi_{\lambda}(\partial_{s}w_{n}(s_{0},\cdot))\to\pi_{\lambda}(\partial_{s}u_{+}(s_{0},\cdot))$ in $C^{\infty}(\mathbb{R}/\mathbb{Z})$ and the latter does not vanish, we get that $\zeta_{n}(s_{0},\cdot)$ does not vanish and $\operatorname{wind}(\zeta_{n}(s_{0},\cdot))=\operatorname{wind}_{\infty}(\widetilde{u}_{+},\infty,\tau)$ for all $n$ large enough. The argument principle now implies that
@@
With (2.19) established, we now invoke the dynamical convexity of $\lambda$ up to action $C$. Namely, at each negative puncture $z\in\Gamma$ we have $\operatorname{wind}_{\infty}(\widetilde{u},z,\tau)\geq 2$ since the corresponding asymptotic limits satisfy ${\rm CZ}^{0}_{\tau}\geq 3$. Hence
@@
Now the exact same argument as above will show that $\operatorname{wind}_{\infty}(\widetilde{v})\leq 1$, i.e. $\widetilde{v}$ is fast. Since $\gamma$ is simply covered we know that $\widetilde{v}$ is somewhere injective. Results from [MW] now imply that $\widetilde{v}$ has to be an embedding, since critical points or self-intersections of $\widetilde{v}$ would force that $\widetilde{v}_{n}$ is not an embedding for $n$ large enough, a contradiction. This concludes the proof that $\mathcal{Y}_{L}$ is compact.
@@
We now produce the Reeb orbit $\gamma$ claimed by the Lemma 2.1, which will be denoted by $\gamma_{-}$ below.
The Fredholm theory from Step 2 provides families of planes in $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ by exponentiating nonvanishing sections of the normal bundle. Hence the set of points in $\overline{W}$ contained in the image of some plane in $\mathcal{Y}$ is open. This fact and the compactness of $\mathcal{Y}_{L}$ proved above together show that we can find a sequence of planes $[\widetilde{v}_{n}]\in\mathcal{Y}$ satisfying
@@
By SFT compactness again, we can pass to a subsequence so that $[\widetilde{v}_{n}]$ converges to a building ${\bf u}$. Unlike in Step 3, ${\bf u}$ now has at least one lower level in (×$(\mathbb{R}\times Y^{-},\widetilde{J})$, in addition to a unique level in $(\overline{W},\bar{J})$, and possibly some upper levels in (×$(\mathbb{R}\times Y^{+},\lambda)$.
The same argument as in Step 3, using (i) and (ii), shows that there are no upper levels in (×$(\mathbb{R}\times Y^{+},\lambda)$. It follows that the level in $(\overline{W},\bar{J})$ consists precisely of a finite energy sphere $\widetilde{u}_{0}:(\mathbb{C}\setminus\Gamma_{0},i)\to(\overline{W},\bar{J})$ with precisely one positive puncture (at $\infty$) where the asymptotic limit is $\gamma_{+}$. If the asymptotic behaviour at the positive puncture is trivial then $\operatorname{wind}_{\infty}(\widetilde{u}_{0},\tau,\infty)=-\infty$. If the the asymptotic behaviour at $\infty$ is nontrivial trivial then, up to reparametrisation, the SFT compactness theorem provides property (iii), and one can argue similarly to Step 3 to prove that $\operatorname{wind}_{\infty}(\widetilde{u}_{0},\tau,\infty)\leq 1$.
We claim now that $\widetilde{u}_{0}$ is an embedded cylinder, asymptotic at its negative puncture to a Reeb orbit $\gamma_{-}$ satisfying ${\rm CZ}(\gamma_{-})=3$. To see this, first note that $\widetilde{u}_{0}$ is somewhere injectiv,e because the asymptotic limit at the unique positive puncture $\infty$ is the simple Reeb orbit $\gamma_{+}$. Since it is a limit of embeddings, the results from [MW] imply that $\widetilde{u}_{0}$ is also an embedding. Now one builds a weighted Fredholm theory with sections of the normal bundle for embedded curves in $(\overline{W},\bar{J})$ with the same asymptotic behaviour as $\widetilde{u}_{0}$. There is only the nontrivial weight $\delta<0$ at the positive where the asymptotic limit is $\gamma_{+}$. The weighted Fredholm index is
${\rm CZ}^{\delta}_{\tau}(\gamma_{+})-\sum_{z\in\Gamma_{0}}{\rm CZ}^{0}_{\tau}(\gamma_{z})-1+\#\Gamma_{0}$
Here $\gamma_{z}$ denotes the asymptotic limit of $\widetilde{u}_{0}$ at the negative punctures $z\in\Gamma_{0}$. Standard arguments imply that if $\bar{J}$ is chosen in a residual subset of almost complex structures, then this the weighted Fredholm index is nonnegative. Assume that $\bar{J}$ has been chosen this way in Step 1. We know that ${\rm CZ}^{\delta}_{\tau}(\gamma)=3$ and ${\rm CZ}^{0}_{\tau}(\gamma_{z})\geq 3$ by dynamical convexity. Plugging this in to the above formula for the weighted Fredholm index, we obtain
@@
It follows that #$\#\Gamma_{0}=1$ and the above inequalities are equalities. Thus $\widetilde{u}_{0}$ is an embedded cylinder asymptotic at its negative puncture to a Reeb orbit $\gamma_{-}$ satisfying ${\rm CZ}(\gamma_{-})=3$.
@@
In [fast, hryn_jsg] it is proved444 In the light of results from [sief_int] this can now be explained in more modern language. If the embedded fast plane $\widetilde{u}=(a,u)$ in (×$(\mathbb{R}\times\partial W,\widetilde{J})$ has as asymptotic limit a simply covered periodic orbit $\hat{P}$ then one can define a weighted self-intersection number $\widetilde{u}*_{\delta}\widetilde{u}={\rm int}(\widetilde{u},\widetilde{u}_{\tau})+\alpha^{<\delta}_{\tau}(\hat{P})$. Here $\tau$ is any $d\lambda$-symplectic trivialization of $\xi$ along $\hat{P}$, and $\widetilde{u}_{\tau}=(a,u_{\tau})$ where $u_{\tau}$ is the map $u$ pushed in the direction of $\tau$ near the puncture. Arguing as in [sief_int], it follows that $\widetilde{u}*_{\delta}\widetilde{u}\geq 0$ and that if $\widetilde{u}*_{\delta}\widetilde{u}=0$ then $u$ is an embedding into $\partial W\setminus\hat{P}$. Plugging $\tau=\tau_{N}$, where $\tau_{N}$ is the trivialization that extends over the normal bundle, we get ${\rm int}(\widetilde{u},\widetilde{u}_{\tau_{N}})=0$ and $\alpha^{<\delta}_{\tau}(\hat{P})=\alpha^{<\delta}_{\tau}(\hat{P})-1=1-1=0$. Hence $\widetilde{u}*_{\delta}\widetilde{u}=0$ and $\hat{P}$ is unknotted. Since $u$ is transverse to the Reeb vector field, the self-linking number of $\hat{P}$ is $-1$. that a simple Reeb orbit that is the asymptotic limit of an embedded fast finite energy plane is unknotted and has self-linking number $-1$. Thus $\gamma_{-}$ is a Hopf orbit, and this completes the proof of Lemma 2.1.

### v3

@@ 2.5. The main compactness argument
Fix $\delta<0$ in the interval between eigenvalues of the asymptotic operator $L_{\gamma_{+},\widetilde{J}}$ satisfying $\operatorname{wind}_{\tau}=1$ and $\operatorname{wind}_{\tau}=2$. This is possible since it follows from dynamical convexity that $\operatorname{CZ}^{0}_{\tau}(\gamma_{+})\geq 3$. Note that $\alpha^{<\delta}_{\tau}(\gamma_{+})=1$ and $\operatorname{CZ}^{\delta}_{\tau}(\gamma_{+})=3$. Following [props3], one can build a Fredholm theory with weight $\delta$ based on sections of the normal bundle. In more detail, let $\widetilde{u}=(a,u)$ represent an arbitrary element of $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$. Using the asymptotic behavior from Theorem 2.5 and the fact that $\widetilde{u}$ is assumed to be an embedding (an immersion would suffice), one can consider a $\bar{J}$-invariant normal bundle $N$ of $\widetilde{u}$ that agrees with $u^{*}\xi$ near the puncture $\infty$. Then one looks at sections of the normal bundle that decay faster then $e^{\delta s}$ near $\infty$, where a neighbourhood of $\infty$ is parametrised by $(s,t)\simeq e^{2\pi(s+it)}$, $s\gg 1$. One can use here Sobolev or Hölder spaces. The linearization of the Cauchy-Riemann equations at $\widetilde{u}$ determines a Fredholm operator $D_{\widetilde{u}}$ on this space of sections with index
${\rm ind}_{\delta}(\widetilde{u})=\operatorname{CZ}^{\delta}_{\tau}(\gamma_{+})-1=3-1=2.$
Lemma 2.10.
The operator $D_{\widetilde{u}}$ is surjective.
Proof.
The contact structure along the plane can be identified, near the puncture, with the normal bundle of the plane. This follows from the asymptotic formula. We proceed assuming this identification is made. Consider $x+iy$ global coordinates on $\mathbb{C}$, and $(s,t)\in\mathbb{R}\times\mathbb{R}/\mathbb{Z}$ coordinates on $\mathbb{C}\setminus\{0\}$ defined via the biholomorphism $(s,t)\simeq e^{2\pi(s+it)}$. A symplectic trivialization of the normal bundle $N$ can be homotoped to a trivialization that extends over a $d\lambda$-symplectic trivialization $\tau_{N}$ of $x(T\cdot)^{*}\xi$. A nowhere vanishing section $\zeta_{N}$ of $x(T\cdot)^{*}\xi$ such that $\operatorname{wind}_{\tau_{N}}(\zeta_{N})=0$ extends to a nowhere vanishing section of $N$ still denoted by $\zeta_{N}$. Hence, $\{d\widetilde{u}\cdot\partial_{x},\zeta_{N}\}$ is a complex frame of $\widetilde{u}^{*}TW$. A nowhere vanishing section $\zeta_{1}$ of $x(T\cdot)^{*}\xi$ such that $\operatorname{wind}_{\tau}(\zeta_{1})=0$ extends to a nowhere vanishing section of $\widetilde{u}^{*}\xi$ still denoted by $\zeta_{1}$. Hence, $\{\partial_{a},\zeta_{1}\}$ is a complex frame of $\widetilde{u}^{*}TW$. The asymptotic behavior tells us that $\partial_{a}$ is roughly $d\widetilde{u}\cdot\partial_{s}$, up to a positive constant, near the puncture. Since $\mathbb{C}$ is contractible,
$0=\operatorname{wind}(d\widetilde{u}\cdot\partial_{s},d\widetilde{u}\cdot\partial_{x})+\operatorname{wind}(\zeta_{1},\zeta_{N})=1+\operatorname{wind}(\zeta_{1},\zeta_{N}).$
This shows that $\tau_{N}$ is “one turn ahead” of $\tau$. Hence, a nowhere vanishing section of $x(T\cdot)^{*}\xi$ with winding number equal to 0 with respect to $\tau_{N}$ will have winding number equal to +1+1 with respect to $\tau$. In particular,
$\alpha^{<\delta}_{\tau_{N}}(\gamma_{+})=\alpha^{<\delta}_{\tau}(\gamma_{+})-1=1-1=0.$
@@
A section $\zeta$ in the kernel of $D_{\widetilde{u}}$ satisfies a Cauchy-Riemann type equation. Hence, either $\zeta\equiv 0$, or its zeros are isolated and count positively to the total algebraic count of zeros. Moreover, if $\zeta\not\equiv 0$ then its asymptotic behavior at $\infty$ is governed by an eigenvector $e(t)$ of the asymptotic operator $L_{\gamma_{+},\widetilde{J}}$ associated to an eigenvalue $\nu<0$ in a similar manner to what is described by Theorem 2.5: $\zeta(s,t)\sim e^{\nu s}(e(t)+\varepsilon(s,t))$, $\sup_{t}|\varepsilon(s,t)|\to 0$ as $s\to+\infty$. One can justify this asymptotic behavior using [sie_CPAM, Theorem A.1]. Hence $\zeta$ has only finitely many zeros, and by the argument principle the total count of zeros of $\zeta$ is nonnegative and equal to $\operatorname{wind}_{\tau_{N}}(\nu)$. The exponential decay $e^{\delta s}|\zeta(s,t)|\to 0$ forces $\nu<\delta$. Thus $\operatorname{wind}_{\tau_{N}}(\nu)\leq\alpha^{<\delta}_{\tau_{N}}(\gamma_{+})=0$, and we conclude that $\zeta$ never vanishes if it does not vanish identically. If $D_{\widetilde{u}}$ is not surjective then its kernel has dimension $\geq 3$ since the index is $2$, so we would find three linearly independent sections in the kernel, but this is impossible: a linear combination of them would vanish at some point because the normal bundle has rank two. This proves the desired surjectivity. ∎
@@
The automatic transversality proved above endows $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ with the structure of a two-dimensional manifold. Note that since the Fredholm theory is modelled on a space of sections with exponential decay faster than $\delta$, the asymptotic eigenvalue of the obtained nearby planes is always $\delta$, implying that $\operatorname{wind}_{\infty}\leq\alpha^{<\delta}_{\tau}=1$ for the nearby planes.
@@
In preparation for compactness arguments below, we note that one can use results on cylinders of small area, such as those proven in [small_area], together with the assumed nondegeneracy of $\gamma_{+}$, to conclude that the topology $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ inherits from the functional analytic set-up agrees with the topology induced by $C^{\infty}_{\rm loc}$-convergence. This is the property of “completeness” in [props3].
@@
Suppose now that $\widetilde{u}$ is an embedded fast plane in $\overline{W}$ asymptotic to $\gamma_{+}$. Let $\mathcal{Y}\subset\mathcal{M}^{\rm fast}(\gamma,\bar{J})$ denote the connected component of $\widetilde{u}$]. For each $L\geq 0$ consider the compact subset of $\overline{W}$ defined by
@@
$\mathcal{Y}_{L}=\{[\widetilde{v}]\in\mathcal{Y}\mid\widetilde{v}(\mathbb{C})\cap E_{L}\neq\emptyset,\ \widetilde{v}(\mathbb{C})\cap\left((-\infty,-L)\times Y^{-}\right)=\emptyset\}.$
@@
We first prove that $\bf u$} has only the bottom level. Suppose by contradiction that $\bf u$} has levels in $\mathbb{R}\times Y^{+}$. Then the top level is a finite energy punctured sphere
[…]
with precisely one positive puncture (at $\infty$) where it is asymptotic to $\gamma_{+}$, and negative punctures at $\Gamma$. If $\int u_{+}^{*}d\lambda=0$, then $\widetilde{u}_{+}$ is the trivial cylinder over $\gamma_{+}$ since $\gamma_{+}$ is simply covered, contradicting the nontriviality in the previous paragraph. Hence $\int u_{+}^{*}d\lambda>0$ and $\widetilde{u}_{+}$ has nontrivial asymptotic behavior at all its punctures.
@@
Note that there is a uniformity in neighbourhoods of $\infty$ encoded in property (i). Write $(s,t)$ instead of $e^{2\pi(s+it)}$. Theorem 2.5 provides a constant $s_{0}>0$ such that $\pi_{\lambda}(\partial_{s}u_{+})$ does not vanish on $s_{0},+\infty)\times\mathbb{R}/\mathbb{Z}$. In particular the winding number $\operatorname{wind}_{\tau}(\pi_{\lambda}(\partial_{s}u_{+}(s_{0},\cdot)))$ of $t\mapsto\pi_{\lambda}(\partial_{s}u_{+}(s_{0},t))$ in the global frame $\tau$ is equal to $\operatorname{wind}_{\infty}(\widetilde{u},\infty,\tau)$. By (i) we can assume, up to making $s_{0}$ larger, that $\pi_{\lambda}(\partial_{s}w_{n})$ is defined on the set $s_{0},+\infty)\times\mathbb{R}/\mathbb{Z}$. Use $\tau$ to represent $\pi_{\lambda}(\partial_{s}w_{n})$ as a $\mathbb{C}$-valued function $\zeta_{n}(s,t)$. This function satisfies a Cauchy-Riemann type equation, hence its zeros count positively towards the total algebraic count of zeros. Since $\pi_{\lambda}(\partial_{s}w_{n}(s_{0},\cdot))\to\pi_{\lambda}(\partial_{s}u_{+}(s_{0},\cdot))$ in $C^{\infty}(\mathbb{R}/\mathbb{Z})$ and the latter does not vanish, we get that $\zeta_{n}(s_{0},\cdot)$ does not vanish and $\operatorname{wind}(\zeta_{n}(s_{0},\cdot))=\operatorname{wind}_{\infty}(\widetilde{u}_{+},\infty,\tau)$ for all $n$ large enough. The argument principle now implies that
@@
With (2.19) established, we now invoke the dynamical convexity of $\lambda$ up to action $C$. Namely, at each negative puncture $z\in\Gamma$ we have $\operatorname{wind}_{\infty}(\widetilde{u},z,\tau)\geq 2$ since the corresponding asymptotic limits satisfy $\operatorname{CZ}^{0}_{\tau}\geq 3$. Hence
@@
Now the exact same argument as above will show that $\operatorname{wind}_{\infty}(\widetilde{v})\leq 1$, i.e. $\widetilde{v}$ is fast. Since $\gamma$ is simply covered we know that $\widetilde{v}$ is somewhere injective. An isolated self-intersection of $\widetilde{v}$ counts strictly positively to the total self-intersection number, forcing self-intersections of $\widetilde{v}_{n}$ for $n$ large enough, a contradiction. Non-isolated self-intersections would force $\widetilde{v}$ to be multiply covered, a contradiction. A direct application of [MW, Theorem 7.2] implies that $\widetilde{v}$ does not have critical points: critical points of $\widetilde{v}$ are isolated, and $\widetilde{v}$ is a pseudo-holomorphic map that is a $C^{\infty}_{\rm loc}$-limit of pseudo-holomorphic maps $\widetilde{v}_{n}$. Hence, $\widetilde{v}$ is an embedding.
In fact, This is a contradiction. Non-isolated self-intersections would force $\widetilde{v}$ to be multiply covered, contradicting the fact that $\gamma_{+}$ is a simple orbit.
Similarly, a critical point of $\widetilde{v}$ would force that $\widetilde{v}_{n}$ is not an embedding for $n$ large enough, a contradiction. This concludes the proof that $\mathcal{Y}_{L}$ is compact.
@@
We now produce the Reeb orbit $\gamma$ claimed by the Lemma 2.1. From now on $\gamma_{+}$ is assumed to be a Hopf orbit with action $\mathcal{A}_{\rm Hopf}(\Omega_{2})$. Existence of $\gamma_{+}$ follows from the existence of Hopf orbits in $Y^{+}$ together with an application of Lemma 2.2.
The moduli space $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ is nonempty. To prove this it suffices to show that there exists an embedded fast finite-energy plane in the symplectization $\mathbb{R}\times Y^{+}$ asymptotic to $\gamma_{+}$. This is proved in [hryn_jsg, Lemma 3.14].
The Fredholm theory from Step 2 provides families of planes in $\mathcal{M}^{\rm fast}(\gamma_{+},\bar{J})$ by exponentiating nonvanishing sections of the normal bundle. Hence the set of points in $\overline{W}$ contained in the image of some plane in $\mathcal{Y}$ is open. This fact and the compactness of $\mathcal{Y}_{L}$ proved above together show that we can find a sequence of planes $\widetilde{v}_{n}]\in\mathcal{Y}$ satisfying
@@
[…]
The same argument as in Step 3, using (i) and (ii), shows that there are no upper levels in $(\mathbb{R}\times Y^{+},\lambda)$. It follows that the level in $(\overline{W},\bar{J})$ consists precisely of a finite energy sphere $\widetilde{u}_{0}:(\mathbb{C}\setminus\Gamma_{0},i)\to(\overline{W},\bar{J})$ with precisely one positive puncture (at $\infty$) where the asymptotic limit is $\gamma_{+}$. If the asymptotic behavior at the positive puncture is trivial then $\operatorname{wind}_{\infty}(\widetilde{u}_{0},\tau,\infty)=$-$\infty$. If the asymptotic behavior at $\infty$ is nontrivial trivial then, up to reparametrisation, the SFT compactness theorem provides property (iii), and one can argue similarly to Step 3 to prove that $\operatorname{wind}_{\infty}(\widetilde{u}_{0},\tau,\infty)\leq 1$.
We claim now that $\widetilde{u}_{0}$ is an embedded cylinder, asymptotic at its negative puncture to a Reeb orbit $\gamma_{-}$ satisfying $\operatorname{CZ}(\gamma_{-})=3$. To see this, first note that $\widetilde{u}_{0}$ is somewhere injective because the asymptotic limit at the unique positive puncture $\infty$ is the simple Reeb orbit $\gamma_{+}$. Since it is a limit of embeddings, the results from [MW] imply that $\widetilde{u}_{0}$ is also an embedding. Now one builds a weighted Fredholm theory with sections of the normal bundle for embedded curves in $(\overline{W},\bar{J})$ with the same asymptotic behavior as $\widetilde{u}_{0}$. There is only the nontrivial weight $\delta<0$ at the positive end where the asymptotic limit is $\gamma_{+}$. The weighted Fredholm index is
$\operatorname{CZ}^{\delta}_{\tau}(\gamma_{+})-\sum_{z\in\Gamma_{0}}\operatorname{CZ}^{0}_{\tau}(\gamma_{z})-1+\#\Gamma_{0}$
Here $\gamma_{z}$ denotes the asymptotic limit of $\widetilde{u}_{0}$ at the negative punctures $z\in\Gamma_{0}$. Standard arguments imply that if $\bar{J}$ is chosen in a residual subset of almost complex structures, then this weighted Fredholm index is nonnegative. Assume that $\bar{J}$ has been chosen this way in Step 1. We know that $\operatorname{CZ}^{\delta}_{\tau}(\gamma)=3$ and $\operatorname{CZ}^{0}_{\tau}(\gamma_{z})\geq 3$ by dynamical convexity. Plugging this in to the above formula for the weighted Fredholm index, we obtain
@@
It follows that $\Gamma_{0}=1$ and the above inequalities are equalities. Thus $\widetilde{u}_{0}$ is an embedded cylinder asymptotic at its negative puncture to a Reeb orbit $\gamma_{-}$ satisfying $\operatorname{CZ}(\gamma_{-})=3$.
@@
[…]